### v3

# Connecting the Dots: Evaluating Abstract Reasoning Capabilities of LLMs Using the New York Times Connections Word Game

Abstract
The New York Times Connections game has emerged as a popular and challenging pursuit for word puzzle enthusiasts. We collect 200 Connections games to evaluate the performance of state-of-the-art large language models (LLMs) against expert and novice human players. Our results show that even the best-performing LLM, GPT-4o, which has otherwise shown impressive reasoning abilities on a wide variety of benchmarks, can only fully solve 8% of the games. Compared to GPT-4o, novice and expert players perform better, with expert human players significantly outperforming GPT-4o. We create a taxonomy of the knowledge types required to successfully cluster and categorize words in the Connections game, revealing that LLMs struggle with associative, encyclopedic, and linguistic knowledge. Our findings establish the New York Times Connections game as a challenging benchmark for evaluating abstract reasoning capabilities in humans and AI systems.
Connecting the Dots: Evaluating Abstract Reasoning Capabilities of LLMs Using the New York Times Connections Word Game
Prisha Samadarshi2     Mariam Mustafa2     Anushka Kulkarni2     Raven Rothkopf2      Tuhin Chakrabarty1thanks: *Equal contribution.     Smaranda Muresan1,2* 1 Department of Computer Science, Columbia University 2 Barnard College {ps3203, mm5970, ajk2256, rgr2124}@barnard.edu, tuhin.chakr@cs.columbia.edu, smara@cs.columbia.edu
1 Introduction
Word puzzle enthusiasts have become captivated by Connections, an engaging game launched by the New York Times (NYT) in June 2023. This daily game presents players with a 4x4 grid containing 16 words and tasks them to identifying four distinct clusters that link the corresponding four words in each cluster through some shared characteristics (Figure 1 [a] vs [b]). Despite its seemingly straightforward premise, Connections delivers a stimulating linguistic workout that keeps players returning daily to test their mental acuity and semantic savvy. The categories 1 (yellow), 2 (green), 3 (blue), and 4 (purple) are arranged according to ascending levels of difficulty. Category 1 is the most intuitive, while Category 4 is the hardest. For instance, in Figure 1 (b), the most straightforward category is "Conformists" {Followers, Lemmings, Puppets, Sheep}, while the most challenging category includes {Apartment, Insults, Likes, Shovels} and requires the understanding that a single word (in this case, "digs") can have multiple meanings that differ in etymology or sense, depending on the context.
(a) The unsolved connections game presented to a player
(b) The solved connections game with distinct categories sorted according to levels of difficulty—straightforward (yellow) to tricky (purple)
Figure 1: Example from a NYT Connections game
While the task may appear easy, many words clump easily into categories, acting as red herrings. For instance Likes, Followers, Shares, Insult might be categorized as “Social Media Interactions" at first glance. However, the game is designed to promote orthogonal thinking and pushes players to find unusual ways to group things. To group words across proper categories, as shown in Figure 1 (b), a player must reason with various forms of knowledge spanning from Semantic Knowledge (Conformists) to Encyclopedic Knowledge (U.S. cities).
Abstract reasoning represents a person’s ability to solve problems, identify patterns, and work with logical systems Barrett et al. (2018); Johnson et al. (2021). While the performance of large language models (LLMs) on arithmetic and language-based commonsense reasoning benchmarks has been the subject of recent analyses, it is unclear whether these LLMs possess abstract reasoning capabilities that are often challenging even for humans Xu et al. (2023). Given its nature, we choose the NYT Connections Game as a test bed for investigating the abstract reasoning capabilities of both humans and large language models (LLMs). We collect 200 distinct Connection games and test the capabilities of four state-of-the-art large language models, namely Google’s Gemini 1.5 Pro Team et al. (2023), Anthropic’s Claude 3 Opus Anthropic (2024), OpenAI’s GPT4-Omni OpenAI (2023), and Meta’s Llama 3 70B AI@Meta (2024) and compare them with human performance.
While all LLMs can partially solve some of the games, their performance is far from ideal. Experimental evidence shows that with few-shot and chain-of-thought prompting, even the best-performing LLM, GPT-4o, can only solve 8% of the games completely. We recruit human players at novice and expert levels of proficiency and compare their performance to GPT-4o. Our results show that the Connections game serves as a challenging benchmark for reasoning, with novice players performing only marginally better than GPT-4o. On the contrary, expert players perform significantly better than GPT-4o in solving games perfectly (Section 5). To better understand the challenging nature of this benchmark, we create a taxonomy of knowledge required to group words into their respective categories (Section 3.2). Our analysis shows that while LLMs are good at reasoning that involves some types of semantic knowledge, they struggle with other types of knowledge such as associative, encyclopedic, or muti-word expressions. Our code and data will be made available to the public at https://github.com/mustafamariam/LLM-Connections-Solver.
2 Related Work
Advancements in LLMs have led to a growing interest in exploring their potential to take on intricate and conceptual challenges by generating linguistic sequences. These models have shown potential in gaming by serving as players Wang et al. (2023a); Tsai et al. (2023); Ciolino et al. (2020); Bakhtin et al. (2022); Noever et al. (2020), non-player characters (NPCs) Park et al. (2023); Urbanek et al. (2019), and generating game content Todd et al. (2023); Wang et al. (2023b); Sudhakaran et al. (2024); Ammanabrolu and Riedl (2021).
Recent research has explored applying large language models (LLMs) and other natural language processing (NLP) techniques to solve and generate text-based puzzles. Zhao and Anderson (2023) use LLMs to tackle and create the weekly Sunday Puzzles featured on National Public Radio (NPR). When presented with multiple-choice questions, GPT-3.5 attained an accuracy of up to 50%. However, when asked to generate novel and engaging puzzles, the model encounters challenges. Compared to the Connections game, NPR’s weekly puzzles tend to emphasize character-level word transformations and relatively common references, relying less on encyclopedic, associative, or semantic knowledge. Rozner et al. (2021) examined the potential of using "cryptic crossword" clues as an NLP benchmark. Wallace et al. (2022) propose automatic ways of solving crossword puzzles by generating answer candidates for each crossword clue using neural question answering models and combining loopy belief propagation with local search to find full puzzle solutions. Our work builds upon these efforts by utilizing the NYT Connections puzzle as a means to investigate the abstract reasoning capabilities of state-of-the-art LLMs.
Knowledge Category Words
Encyclopedic
TV SHOWS WITH
HAPPY-SOUNDING NAMES
[‘CHEERS’, ‘EUPHORIA’, ‘FELICITY’, ‘GLEE’]
JACKS [‘BLACK’, ‘FROST’, ‘MA’, ‘SPARROW’]
Semantic Synonym COLLEAGUES [‘ASSOCIATE’, ‘FELLOW’, ‘PARTNER’, ‘PEER’]
Polysemy WHAT A MOLE CAN BE [‘ANIMAL’, ‘BIRTHMARK’, ‘SPY’, ‘UNIT’]
Hypernym PERIOD [‘AGE’, ‘DAY’, ‘ERA’, ‘TIME’]
Associative ORIGIN [‘CRADLE’, ‘FONT’, ‘ROOT’, ‘SOURCE’]
THINGS THAT ARE ORANGE [‘BASKETBALL’, ‘CARROT’, ‘GOLDFISH’, ‘PUMPKIN’]
Linguistic NOUN SUFFIXES [‘DOM’, ‘ION’, ‘NESS’, ‘SHIP’]
SILENT “W" [‘ANSWER’, ‘TWO’, ‘WRIST’, ‘WRONG’]
Multiword Expression ___WOOD [‘DOG’, ‘DRIFT’, ‘HOLLY’, ‘SANDAL’]
Combined CITY HOMOPHONES [‘DELI’, ‘NIECE’, ‘ROAM’, ‘SOUL’]
SOCIAL MEDIA APP ENDINGS [‘BOOK’, ‘GRAM’, ‘IN’, ‘TUBE’]
Table 1: Different types of knowledge required to group words into their respective categories
The word association task Galton (1879) has been used extensively in psychological and linguistic research as a way of measuring connections between words in the mental lexicon. Responses in word association tasks have informed what we know about the structure and organization of semantic memory and the mental lexicon De Deyne and Storms (2008). In this work, we similarly show how one must utilize semantic and associative memories to solve the Connections game.
Chollet (2019) proposed the Abstraction and Reasoning Corpus (ARC), built upon an explicit set of priors designed to be as close as possible to innate human priors and argued that it can be used to measure a human-like form of general fluid intelligence, enabling fair general intelligence comparisons between AI systems and humans. Recently Xu et al. (2023) show that GPT-4 solves only 13/50 of the most straightforward ARC tasks, demonstrating a significant gap in the abstract reasoning capabilities of LLMs. Prior work has also studied abstract reasoning in Neural Networks Barrett et al. (2018) even in the presence of distracting features Zheng et al. (2019). Our work builds upon these and presents the Connections game as a compelling benchmark for abstract reasoning capabilities for LLMs in the presence of distractors.
3 Data
3.1 Collection
To gather the necessary data, we found an archival site consisting of all possible answer choices and their corresponding categorizations. As the NYT does not maintain an archive of Connection puzzles, we resorted to an external, third-party site for data collection. [^1] https://tryhardguides.com/nyt-connections-answers/ Our data spans daily problems from the conception of Connections, June 2023, to March 2024. In total, we gather 203 distinct games, out of which 3 are used for few-shot prompting, while the remaining 200 comprise the dedicated test set.
3.2 Types of Reasoning
Semantic
Knowledge
Associative
Knowledge
Encyclopedic
Knowledge
Mutiword
Expressions
Linguistic
Knowledge
Combined
Knowledge
337 171 153 77 49 13
Table 2: Breakdown of instances of different types of reasoning across 800 categories from 200 Connections games
Investigating the relationship between words offers insights into both the structure of language and the influence of cognition on linguistic tasks Stella et al. (2018). To solve Connections games, players must draw on certain aspects of word knowledge, such as a word’s meaning. To deepen our understanding, we bucket each $<$category, grouping$>$ into the types of knowledge that are primarily required to solve them. Two experts annotate a total of 800 samples coming from 200 games into 6 broader categories. On 8.6% of the 800 samples where they disagree (See examples of disagreement in Appendix B), the experts engaged in discussion Schaekermann et al. (2018); Chen and Zhang (2023); Chen et al. (2019) to arrive at an individual category.
3.2.1 Semantic Knowledge
The majority of instances in the Connections game require possessing knowledge of lexical semantics Cruse (1986), particularly semantic relations such as synonymy (words with the same meaning), hypernymy/hyponymy (relation between a generic terms and its specific instance), and polysemy (many possible meanings for a word). Table 1 shows three examples of groups that use such Semantic Knowledge.
3.2.2 Associative Knowledge
To group words into their respective categories one often needs to think beyond the lexical semantic relations mentioned above. Associative learning Shanks (1995) occurs when an element is taught through association with a separate, seemingly unrelated pre-occurring element. To cluster words using Associative Knowledge, one either needs to focus on the connotative meaning of a word or the shared property that connects several words. For instance, as shown in Table 1, the words Cradle, Root, or Font in their literal sense do not refer to Origin; instead, one needs to rely on their connotative meaning for such a categorization. Similarly, on the surface level, Basketball, Carrot, Goldfish, and Pumpkin are unrelated. However, a shared property that connects them is their orange color.
3.2.3 Encyclopedic Knowledge
We notice that to group certain sets of words into their proper categories, one needs knowledge that spans beyond concepts and relies on entities in the real world found in knowledge bases such as Wikipedia Mihalcea and Csomai (2007). This can be seen in Table 1, where, to bucket the words Black, Frost, Ma, and Sparrow into the category of Jacks, one needs to possess knowledge across various domains: ‘Jack Black’ an American actor, ‘Jack Frost’ a character from English folklore who personifies winter, ‘Jack Ma’ the founder of Alibaba, and ‘Jack Sparrow’ the protagonist of the Pirates of the Caribbean film series. We label this type of knowledge Encyclopedic Knowledge.
3.2.4 Multiword Expressions
Multiword Expressions are complex constructs that interact in various, often untidy ways and represent a broad continuum between non-compositional (or idiomatic) and compositional groups of words Moon (1998). Higher difficulty levels (blue and purple) in the Connections game often require players to recognize that the four words can form a Multiword Expression if combined with an external word. Table 1 shows examples of Multiword Expressions where half of the expressions are given in the form of individual words and the player needs to find the other half to categorize words into the correct group.
3.2.5 Linguistic Knowledge
Linguistic competence Coseriu (1985) is the system of unconscious knowledge that one has when one knows a language. Such competence is often required to classify words into their appropriate categories. Several instances from the Connections game require knowledge of morphology, phonology, or orthography for correct categorization. For example, as shown in Table 1, one needs knowledge about morphology to group Dom, Ion, Ness, and Ship as Noun Suffixes. Similarly, one needs to rely on phonological knowledge about the sound patterns of Answer, Two, Wrist, and Wrong to categorize them as Silent "W".
3.2.6 Combined
Some of the hardest examples in the NYT Connections game require reasoning with multiple types of knowledge. For instance, the example in Table 1 shows that to group Deli, Niece, Roam, and Soul, one requires the knowledge that these words have the same phonological form with the cities Delhi, Nice, Rome, and Seoul. This categorization requires the simultaneous use of Encyclopedic and Linguistic Knowledge. Similarly, to group the words Book, Gram, In, and Tube together one needs to identify that they are essentially parts of closed compounds (Face+Book, Insta+Gram, Linked+In, You+Tube) that also represent popular social media apps. This categorization requires one to use Encyclopedic and Linguistic Knowledge together.
4 Experimental Settings
4.1 LLMs as Game Players
To test the capabilities of large language models in solving the Connections game, we rely on recent advancements in in-context learning and chain-of-thought prompting Wei et al. (2022). We provide 3 complete examples in our few-shot prompt along with rules and common strategies that players must use to solve the game. We also elicit chain-of-thought reasoning Wei et al. (2022) requiring models to explain their groupings and categories chosen. Formulation of the prompt involved trial and error; the first iteration of the prompt included the Connections game instructions provided by the New York Times (Liu, 2023a), and included three demonstrations with gold labels asking the LLM to explain its reasoning in a step-by-step manner Wei et al. (2022). We ran this first prompt with a few games on a development set of 30 games (different from our test set), using the 4 LLMs. After identifying commonalities in the types of errors made by the LLMs while playing the game, we added additional instructions about the game, specified the response format, and included some tips from a NYT article about playing Connections (Aronow and Levine, 2023). The entire prompt is in Appendix A. To ensure consistency and fairness in performance, we prompt 4 LLMs — Gemini 1.5 Pro, Claude 3 Opus, GPT-4o, and Llama 3 70B —with the same input and use the default sampling parameters (temperature and top_p). We use the scoring schema outlined in Section 4.3 to evaluate how all models perform in solving 200 Connections games spanning from June 15, 2023 to January 1, 2024.
4.2 Humans as Game Players
Alongside LLMs, we recruited 17 human evaluators in two subgroups: 12 novice players with little to no prior experience playing Connections and 5 expert or regular Connections players. The novice and the expert evaluators were peers of the authors of the paper who volunteered to participate without any payment.
We designed a human evaluation interface and randomly sampled 100 games from our test set. Appendix E has more information about the interface. The first screen displays a shortened version of the instructions from the LLM’s prompt so as to not overwhelm the human players. To ensure that the humans solve the game in a manner comparable to the LLMs setup, they were given one try to fully solve the game (i.e., make all 4 categorizations).
Playing these games is a significant cognitive burden. As such, each novice human evaluator played around 8-12 distinct games for a total of 100 randomly sampled games out of the 200 games in the test set, and expert participants each played 10 games for a total of 50 games from the subset of 100 games played by novices.
4.3 Evaluation Criteria
Our scoring schema was developed as a means to numerically interpret the outcome of each Connections game and standardize comparison across LLMs and human players. We outline two processes to obtain clustering and categorical reasoning scores of a game of Connections.
Figure 2: Frequency of unweighted clustering scores for 4 LLMs across 200 games. The number of games in which the respective unweighted clustering score was achieved is atop each bar.
4.3.1 Clustering Score
The clustering score evaluates the ability to correctly group together all the words in the game. We consider two clustering scores. The first, or the unweighted clustering score, is calculated independently of the categories’ supposed difficulty. In this simple scoring mechanism, we allocate one point for each correct cluster (when all 4 words in the group classified by the LLM or human correspond to the 4 words in the gold category/grouping). Ideally, a player’s score should be close to the maximum of 4, signifying that all 4 groups were correctly identified. The equation is as follows:
$score=n_{0}+\ldots+n_{3}$ (1)
where $n_{x}=1$ for each correct grouping $x$ and $n_{x}=0$ for each incorrect grouping.
The second score, referred to as the weighted clustering score, takes into account the difficulty of each grouping. The worst weighted clustering score a player can obtain is 0, meaning that no words were grouped correctly. Ideally, a player’s score should be close to the maximum of 10, signifying that all 4 categories were correctly classified. The equation for this score is as follows:
$score=n_{0}\cdot w_{0}+\ldots+n_{3}\cdot w_{3}$ (2)
where $n_{x}$ represents one of the 4 categories and is always equal to 1 for each category $x$. The reward procedures are as follows: $w_{0}=1$ for a Yellow (most straightforward) correct grouping, $w_{1}=2$ for a Green correct grouping, $w_{2}=3$ for a Blue correct grouping, $w_{3}=4$ for a Purple (trickiest) correct grouping. Our schema for the clustering scores does not incorporate the number of tries as a variable, since in our setup the LLMs are prompted once and take one try to solve the game.
4.3.2 Categorical Reasoning
While the weighted and unweighted clustering scores are calculated for LLMs and humans, the categorical reasoning score is used only for the LLMs’ responses. If all 4 words in a category are correctly identified by an LLM, we conduct further analysis to evaluate whether the LLM reasoned correctly why the words in the groups belong together. We make this distinction in our evaluation so that — in conjunction with the taxonomy knowledge for Connections categories (Section 3.2) — we can assess the types of reasoning that the LLMs are most or least adept in. Since our prompt asks the LLM to include the category name and share the reasons why it grouped words together, we can evaluate whether the LLM’s reasoning in its response is semantically analogous to the gold NYT Connections-provided category name. The decision of semantic equivalence between LLMs output and gold is done manually by a human judge to ensure accuracy.
5 Results
5.1 LLM performance
Overall, we find that GPT-4o performs best across all 200 games. Figure  2 shows the unweighted clustering scores for all 4 LLMs. GPT-4o has the lowest percentage of games in which it made no correct clustering (53 out of 200) and the most games solved perfectly (16 out of 200). Claude 3 Opus is a close second for perfectly solved games at 13 out of 200. Gemini 1.5 Pro performs the worst overall. While it could not make any correct clusters for half of the games, it was able to solve 9 games perfectly, outperforming Llama 3 70B’s 6 games solved perfectly.
Figure 3: Spread of weighted clustering scores for each model across 200 Connections games
In terms of weighted clustering scores for each model (Figure 3), Gemini 1.5 Pro and Llama 3 70B show similar results. Most of their scores are concentrated before 2, showing that these models showed a higher ability to correctly classify the easiest or second easiest categories. GPT-4o and Claude 3 Opus also showed similar results, with most of their weighted clustering scores concentrated before 4, meaning they were better at classifying more and harder categories. Weighted clustering scores $\geq 8$ are very rarely represented in all the models. Appendix D.2 contains a more detailed breakdown.
5.2 Human Performance
In human performance, we measure both novice and expert players against the best overall performing GPT-4o. For the 100 games played by novices and 50 games played by experts, we compare the same 100 and 50 games played by GPT-4o.
5.2.1 Novice Players
Figure 4: Frequency of clustering scores of GPT-4o and 12 novice humans across 100 games
In the 100 games that the novice players completed, their average unweighted clustering score was 1.38, marginally better than GPT-4o’s average of 1.17 on the same 100 games. GPT-4o and novice humans also had similar weighted clustering score distributions. More details are in Appendix D.1. Due to the setup of the human interface, humans could not receive a clustering score of 3 (if humans correctly solve 3 groupings, the 4th is also correct). Because of GPT-4o’s imperfect instruction-following abilities (repeating or omitting a word), it was still able to obtain a clustering score of 3, as shown in Figure 4.
5.2.2 Expert Players
Figure 5: Frequency of clustering scores of GPT-4o and 5 expert humans across 50 games
Expert human players performed significantly better than novices and GPT-4o, with an average clustering score of 3 compared to GPT-4o’s 1.22 (on the same 50 games) and an average weighted clustering score of 7.4 compared to GPT-4o’s 2.32. The distribution of weighted clustering scores is also far more right-skewed (see Appendix D.1 for more). Figure 5 shows that experts perfectly solve over 60% of the 50 games, while GPT-4o only fully solved 5% of the games.
6 Discussion
6.1 What type of reasoning is hardest for LLMs and humans?
To answer this question we rely on our taxonomy of reasoning types we introduced in Section 3.2. The breakdowns of the reasoning types for the 800 categories in our 200-game dataset are shown in Table  2. The patterns in performance across the types of reasoning parallel the LLMs’ overall performance for the most part, with Claude 3 Opus’s performance in Multiword Expressions defying this pattern. The performance in reasoning categories across models is ranked from best to worst as follows: $Semantic>Associative>Encyclopedic>Linguistic>Multiword>Combined$. The model’s performance in these categories corresponds to the frequency with which each category appears, except the category of Multiword Expressions. Only Claude 3 Opus has a greater than zero success rate (4 out of 77) in Multiword Expressions. In Connections games with Multiword Expressions, they are usually the purple category (most difficult), while semantic and associative reasoning appear most often as yellow or green (easier) categories. For Combined categories, no models output correct clusters.
We find that both novice and expert human players are better at all types of reasoning compared to GPT-4o, although neither cluster any Combined Knowledge categories successfully. Novice performance in reasoning categories across models is ranked from best to worst as follows: $Semantic\approx Linguistic>Associative>Encyclopedic>Multiword>Combined$. The greatest difference between GPT-4o and novices is in Multiword Expressions (difference of around 20%) and Linguistic Knowledge (difference of 25%). Expert performance ranked from best to worst: $Semantic>Encyclopedic>Linguistic>Associative>Multiword>Combined$.
For both LLMs and humans, the Combined Knowledge category seems hardest to grasp. Though experts rank low in Multiword Expressions compared with other types of reasoning, they still perform quite highly, with over 60% of Multiword Expression categories grouped correctly. However, perhaps because of a lack of familiarity with the game and types of categories, novice players, like LLMs, struggle with Multiword Expression, achieving an accuracy of just over 20%. Further breakdowns are in Appendix D.
6.2 How much do distractors prevent both LLMs and humans from correct categorization?
The Connections game is often formulated with item overlap in mind, according to the Connections puzzle creator (Liu, 2023b). These distractors, or red herrings, make the game far more challenging. Red herrings can appear in two ways — as a red herring category or red herring word. In the former instance, 3 ultimately unconnected words seem to form a category of their own with 1 word missing. In the latter, a category seems applicable to more than 4 words, but the extras belong to a separate grouping. Examples of each of these types of red herrings are in Appendix C.
Mistakes resulting from red herrings often occur in categories related to Associative Knowledge. Though the words may be associated in one dimension, the LLMs fail to conduct step-by-step reasoning to find another, perhaps more obscure, grouping (in the case of red herring categories) or the outlier (in the case of red herring words).
6.3 How often do LLMs group the words correctly but present incorrect reasons?
To measure the disparity between LLMs making correct clustering and providing the correct reasoning or category name for their choice, we use a measure calculated from the clustering and categorical reasoning scores. Since the categorical reasoning score is the number of categories reasoned correctly and the clustering score considers whether the grouping was correct independent of the reason behind it, unweighted clusteringcategorical $\frac{\text{unweighted clustering}}{\text{categorical reasoning}}$ tells us how common it is for LLMs to cluster categories correctly by chance.
Model Average Ratio
Gemini 1.5 Pro 0.78
Claude 3 Opus 0.87
GPT-4o 0.86
Llama 0.76
Table 3: Average categorical reasoning to unweighted clustering score ratio by model
The average ratios in Table 3 are fairly high, close to or above 80%. The highest overall performing models GPT-4o and Claude 3 Opus have the highest ratios as well. Though it is fairly uncommon that a model will correctly group without correctly reasoning, there are very few instances where models received both a clustering score of 4 (fully solved game) and a categorical reasoning score of 4.
6.4 How can future work improve on such a benchmark?
Certain strategies grounded in reflective thinking could improve performance on such a benchmark. Instead of greedily choosing the first grouping, identifying red herring words or categories first could prevent the possibility of misclassification. Allowing LLMs to solve the game one category at a time and incorporating the feedback present to humans in the NYT Connections game, including whether a grouping is correct (and what difficulty level it is by color), incorrect, or one word away from a correct grouping, may improve performance as well. Retrieval Augmentation from WordNet or dictionaries for lexical connotations Allaway and McKeown (2020) could further improve such categorization. Finally, creating synthetic training data and training an LLM on this task could further close the gap between expert human and LLM performance. We leave such exploration for future work.
7 Conclusions
We introduced NYT Connection games as a benchmark to test abstract reasoning in state-of-the-art LLMs and evaluate their performance against expert and novice human players. We find that GPT-4o performs best, although it is still no match for expert human players. By examining the performance through our knowledge taxonomy, we obtain a more solid understanding of areas in which LLMs can improve to solve clustering scores. They are fairly deficient in certain types of reasoning required to be a skilled Connections player. Although most possess adequate semantic and associative reasoning capabilities, they struggle with Multiword Expressions and Combined Knowledge categories. Additional struggles arise because they cannot identify red herrings and use step-by-step logic to work around them. Ultimately, we find that excelling in Connections means having a breadth of different knowledge types, and LLMs are unfortunately not yet suited for the task.
8 Limitations
Many of the limitations in this section stem from the lack of data available for Connections games and disparities in the comparison between LLMs and humans. Because it is a fairly recent invention and only one puzzle is released per day, there are only a few hundred games available. Since there are some category patterns learned through frequent play, ideally, a model trained on past Connections games might bridge the gap between LLM are expert human evaluators performance.
We acknowledge that human evaluators were not required to add justifications for the groupings they made. This could have made performance comparisons between humans and LLMs for the types of reasoning more equal. Additionally, because a score of 3 was impossible in the human evaluation interface, we cannot be certain that humans were adept in the type of knowledge of their last category grouped, as this could simply been a matter of grouping all options left. Other limitations of human evaluators include that because they were all peers or acquaintances of the paper’s authors, sampling bias could exist. Though the age range of the humans recruited was 14-60, other demographic factors that may not have been accounted for in this sample.
9 Ethical Considerations
We collect the names of users in the human evaluation game’s database simply for logistical purposes. Other than this, no personal data is collected. The data collected and its purpose were verbally conveyed to each evaluator before asking for their consent. We remove the names of evaluators in the data release. Besides this, we ensure that now and in the future, any data collection is transparent with users and is used in an ethical and responsible manner. Since our research primarily evaluates reasoning in a game environment, there are fewer potential real-world risks of its applications. However, biases in LLMs may be reproduced.
Appendix A Prompt
Solve today’s NYT Connections game. Here are the instructions for how to play this game:
Find groups of four items that share something in common.
Category Examples:
FISH: Bass, Flounder, Salmon, Trout
FIRE ___: Ant, Drill, Island, Opal
Categories will always be more specific than ’5-LETTER-WORDS’, ‘NAMES’, or ’VERBS.’
Example 1:
Words
: [’DART’, ‘HEM’, ‘PLEAT’, ‘SEAM’, ‘CAN’, ‘CURE’, ‘DRY’, ‘FREEZE’, ‘BITE’, ‘EDGE’, ‘PUNCH’, ‘SPICE’, ‘CONDO’, ‘HAW’, ‘HERO’, ‘LOO’]
Groupings:
  1. 1.
    Things to sew: [‘DART’, ‘HEM’, ‘PLEAT’, ‘SEAM’]
  2. 2.
    Ways to preserve food: [CAN’, ‘CURE’, ‘DRY’, ‘FREEZE’]
  3. 3.
    Sharp quality: [’BITE’, ‘EDGE’, ‘PUNCH’, ‘SPICE’]
  4. 4.
    Birds minus last letter: [’CONDO’, ‘HAW’, ‘HERO’, ‘LOO’]
Example 2:
Words
: [’COLLECTIVE’, ‘COMMON’, ‘JOINT’, ‘MUTUAL’, ‘CLEAR’, ‘DRAIN’, ‘EMPTY’, ‘FLUSH’, ‘CIGARETTE’, ‘PENCIL’, ‘TICKET’, ‘TOE’, ‘AMERICAN’, ‘FEVER’, ‘LUCID’, ‘PIPE’]
Groupings:
  1. 1.
    Shared: [’COLLECTIVE’, ‘COMMON’, ‘JOINT’, ‘MUTUAL’]
  2. 2.
    Rid of contents: [’CLEAR’, ‘DRAIN’, ‘EMPTY’, ‘FLUSH’]
  3. 3.
    Associated with “stub”: [’CIGARETTE’, ‘PENCIL’, ‘TICKET’, ‘TOE’]
  4. 4.
    __ Dream: [ ’AMERICAN’, ‘FEVER’, ‘LUCID’, ‘PIPE’])
Example 3:
Words
: [’HANGAR’, ‘RUNWAY’, ‘TARMAC’, ‘TERMINAL’, ‘ACTION’, ‘CLAIM’, ‘COMPLAINT’, ‘LAWSUIT’, ‘BEANBAG’, ‘CLUB’, ‘RING’, ‘TORCH’, ‘FOXGLOVE’, ‘GUMSHOE’, ‘TURNCOAT’, ‘WINDSOCK’]
Groupings:
  1. 1.
    Parts of an airport: [’HANGAR’, ‘RUNWAY’, ‘TARMAC’, ‘TERMINAL’]
  2. 2.
    Legal terms: [‘ACTION’, ‘CLAIM’, ‘COMPLAINT’, ‘LAWSUIT’]
  3. 3.
    Things a juggler juggles: [’BEANBAG’, ‘CLUB’, ‘RING’, ‘TORCH’]
  4. 4.
    Words ending in clothing: [’FOXGLOVE’, ‘GUMSHOE’, ‘TURNCOAT’, ‘WINDSOCK’]
Categories share commonalities:
  • There are 4 categories of 4 words each
  • Every word will be in only 1 category
  • One word will never be in two categories
  • As the category number increases, the connections between the words and their category become more obscure. Category 1 is the most easy and intuitive and Category 4 is the hardest
  • There may be a red herrings (words that seems to belong together but actually are in separate categories)
  • Category 4 often contains compound words with a common prefix or suffix word
  • A few other common categories include word and letter patterns, pop culture clues (such as music and movie titles) and fill-in-the-blank phrases
You will be given a new example (Example 4) with today’s list of words. First explain your reason for each category and then give your final answer following the structure below (Replace Category 1, 2, 3, 4 with their names instead)
Groupings:
Category1: [word1, word2, word3, word4]
Category2: [word5, word6, word7, word8]
Category3: [word9, word10, word11, word12]
Category4: [word13, word14, word15, word16]
Remember that the same word cannot be repeated across multiple categories, and you need to output 4 categories with 4 distinct words each. Also do not make up words not in the list. This is the most important rule. Please obey
Example 4:
Words : [InsertGame]
Groupings
Appendix B Disagreements in Annotations
There was certain extent of disagreements between Semantic and Encyclopedic Knowledge. For instance [’PIKE’, ’SPLIT’, ’STRADDLE’, ’TUCK’] are GYMNASTICS POSITIONS and requires domain specific knowledge so could be thought of as Encyclopedic knowledge but it could be classified under Semantic Knowledge (Type Of relation) as many of these words appear in Wordnet. Some disagreements also occurred between between Associative and Encyclopedic Knowledge. For instance here the shared property for [’BASE’, ’BOND’, ’ELEMENT’, ’SOLUTION’] being CHEMISTRY TERMS requires using Associative Knowledge but this could still require Encyclopedic Knowledge about Chemistry. However since we consider Encyclopedic knowledge only as ones related to a Knowledge Base and entities instead of domain concepts we treat this as Associative Knowledge. Finally due to their colloquial use in the English language sometimes there can be confusion amongst what Semantic and Associative Knowledge. For instance [’BOMB’, ’DUD’, ’FLOP’, ’LEMON’] can be thought of synonyms of FAILURE and hence fall under category of Semantic Knowledge , however Lemon is rarely used for Failure and requires using connotative knowledge (shared property) and hence falls more appropriately under Associative Knowledge.
Appendix C Red Herrings
Figure 6: Example of red herring category where the 3 words outlined in red might seem as though they belong together.
In the puzzle in Figure 6, a red herring category is present. In the highest performing models Claude 3 Opus and GPT-4o created a category called "Milk" with Whole, Skim, and Soy and included a random fourth word that did not fit. Each of these three words, however, belongs to a different category: Whole to Kinds of Numbers, Skim to Touch Lightly, and Soy to Sauces in Chinese Cuisine. In other puzzles including a red herring category like this one, all models make similar rationalizations.
Figure 7: Example of red herring word where the 5 words outlined in red may seem like they belong together.
The game in Figure 7 is an example of a game with a red herring word. The five words that appear as though they belong together are outlined in red. However, Mistletoe, Reindeer, Snowman, and Stocking form the Christmas Related category, while Candy Cane belongs to the category Things with Stripes. In this game, GPT-4o, Gemini 1.5 Pro, and Llama 3 70B all made the mistake of grouping Candy Cane with some combination of three of the other Christmas-related words.
Appendix D Performance
D.1 Humans
The frequency of clustering scores for novice human players in 100 games and scores for GPT-4o in the same games are shown in Table 4. The frequency of clustering scores for expert human players in 50 games and scores for GPT-4o in the same games are in Table 5.
Figures 8 and 9 show the distribution for the weighted clustering scores of GPT-4o against novice humans and expert humans, respectively. Finally, Figures 10 and 11 depict the performance of novice and expert humans, respectively, by reasoning type from our taxonomy of knowledge. Because the total counts of types of reasoning required across the 400 (novice) and 200 (expert) categories are unbalanced, the count of the categories reasoned correctly is shown above each bar.
Unweighted
Clustering
Score
GPT-4o Novice Humans
0 31 30
1 35 39
2 29 13
3 2 0
4 8 18
Table 4: Frequency of clustering scores 0-4 for GPT-4o and novice human players across 100 Connections games
Unweighted
Clustering
Score
GPT-4o Expert Humans
0 16 2
1 18 7
2 10 9
3 1 0
4 5 32
Table 5: Frequency of clustering scores 0-4 for GPT-4o and expert human players across 100 Connections games
Figure 8: Spread of weighted clustering score for GPT-4o and novice human players across 100 Connections games
Figure 9: Spread of weighted clustering score for GPT-4o and expert human players across 50 Connections games
Figure 10: Percentage of categories from each knowledge type correctly classified and reasoned by GPT-4o and novice human players across 100 games. The counts of categories correctly reasoned are displayed above each bar.
Figure 11: Percentage of categories from each knowledge type correctly classified and reasoned by GPT-4o and expert human players across 50 games. The counts of categories correctly reasoned are displayed above each bar.
D.2 LLMs
Table 6 shows the frequency of the unweighted clustering scores (number of categories correctly grouped) for each LLM. The total number of games played by each model is 200. Table 7 is slightly different and shows the frequency of categorical reasoning scores (the categories correctly grouped and reasoned) for each model. Because a caveat for receiving a categorical reasoning score greater than 0 is matching gold categories, a score of 0 is more common than in the unweighted clustering scores.
Figure 12 shows the performance of each LLM by reasoning type from our taxonomy of knowledge. Because the total counts of types of reasoning required across the 800 categories are unbalanced, the count of the categories reasoned correctly is shown above each bar.
Unweighted
Clustering
Score
Gemini 1.5 Pro Claude 3 Opus GPT-4o Llama 3 70B
0 100 68 53 84
1 64 72 77 73
2 26 37 48 32
3 1 10 6 5
4 9 13 16 6
Table 6: Frequency of unweighted clustering scores 0-4 for 4 LLMs across 200 Connections games
Categorical
Reasoning Score
Gemini 1.5 Pro Claude 3 Opus GPT-4o Llama 3 70B
0 116 80 59 98
1 59 66 87 73
2 18 36 41 23
3 7 13 11 4
4 0 5 2 2
Table 7: Frequency of categorical reasoning scores 0-4 for 4 LLMs across 200 Connections games
Figure 12: Percentage of categories from each knowledge type correctly classified and reasoned by the models across 200 games. The counts of categories correctly reasoned are displayed above each bar.
Appendix E Human Evaluation Interface
Figure 13 shows the two main screens of the evaluation interface provided to both novice and expert human evaluators. (a) is the instruction screen, while (b) is an example of a game screen after the user hits the "Play" button. To solve the game in one shot, all 16 words from a game are displayed on the screen in separate boxes, with one drop-down per box. The drop-down consists of four labels: Group 1, Group 2, Group 3, and Group 4. The user’s job is to create 4 groups of 4 words using the given labels. Because the groups are chosen from a drop-down menu where the default option is Group 1, a clustering score of 3 is impossible.
We stored the data collected in a SQLite database. Other than any name of choice users were prompted to enter in the "Name" text entry box, no personal data was collected. Each evaluator was then assigned initials in the final dataset collected for evaluation. These initials are not included in the data release. The data that would be collected and its purpose were verbally conveyed to each evaluator before asking for their consent.
(a) Instruction screen
(b) Example of game play
Figure 13: Human evaluation interface